OpenStudy (anonymous):

Use the Pythagorean Theorem to find the length of the missing side of the right triangle with both legs equal to √2.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

A. 2 B. 2 √2 C. √5 D. None of the above

OpenStudy (misty1212):

\(\sqrt2^2+\sqrt2^2=2+2=4\) and \(\sqrt4=2\)

OpenStudy (anonymous):

so a?

OpenStudy (anonymous):

@misty1212

OpenStudy (misty1212):

yes